OpenStudy (anonymous):

At greater depths, a diver uses the air in the tank more quickly. Assume that the amount of time the air will last is inversely proportional to the pressure at the depth of the dive. • Suppose a tank has enough air to last 60 min at the surface. How long will it last at a depth of 99 ft? (The pressure is 4 atm, or 4 times as great.) • Make a table showing how long the air will last at 0 ft, 20 ft, 33 ft, 40 ft, 50 ft, 66 ft and 99 ft.

OpenStudy (anonymous):

t = k*(1/p) where k is a constant. k is evaluated as follows on the surface of water: 60 = k*(1/1atm) = k -> k = 60atm*min At a depth of 99ft, the pressure is 4atm, so p=4: t = k*(1/p) = (60)*(1/4)

OpenStudy (anonymous):

Alright thank you!

OpenStudy (anonymous):

Np.